OpenStudy (anonymous):

A vending machine offers 8 different drinks. One day, 6 employees each purchased a drink from the vending machine. Find the probability that at least 2 employees purchased the same drink. Round your answer to the nearest hundredth.

OpenStudy (anonymous):

for the 1st one. total cases: each man has 8 choices.so total cases = 8^6. each has different drink: cases = 8*7*6*5*4*3. for at least two have same drink = total cases - when each have different drink = 8^6 - 8*7*6*5*4*3. probability = (8^6 - 8*7*6*5*4*3) / (8^6).